OpenStudy (anonymous):

Find the period of the function. y = -3 cos1/4x Things I know: f(x)=a cosbx then period=2pi/b So: -3cos1/4x period=2pi/(1/4x) I got that but it's not matching my choices?: -3 3pi/4 8π pi/4

OpenStudy (mertsj):

Memorize this: the coefficient of x = 2pi divided by the period or:

OpenStudy (mertsj):

\[c=\frac{2\pi}{period} \] \[period = \frac{2\pi}{c}=\frac{2\pi}{\frac{1}{4}}=2\pi \times \times 4=8\pi \]

OpenStudy (anonymous):

Oh I see. Thanks!

OpenStudy (mertsj):

yw